OpenStudy (anonymous):

what is special about 1729

OpenStudy (amistre64):

is it something to do with primes?

OpenStudy (curry):

1729 is known as the Hardy–Ramanujan number after a famous anecdote of the British mathematician G. H. Hardy regarding a hospital visit to the Indian mathematician Srinivasa Ramanujan. In Hardy's words:[1] “ I remember once going to see him when he was ill at Putney. I had ridden in taxi cab number 1729 and remarked that the number seemed to me rather a dull one, and that I hoped it was not an unfavorable omen. "No," he replied, "it is a very interesting number; it is the smallest number expressible as the sum of two cubes in two different ways." ” The two different ways are these: 1729 = 13 + 123 = 93 + 103